OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol.

OpenStudy (anonymous):

@mathman806

OpenStudy (anonymous):

C=.765+.06(76.5) C=.765+4.95 C=5.355 100-5.355 the total cost is 94.66

OpenStudy (anonymous):

my bad messed up C=.765(100)+.06(.765*100) C=76.5+.06(76.5) C=76.5+4.59 C=81.09 Chris pays 81.09